OpenStudy (anonymous):

Medal and fan awarded which expressioon is the simplified form of sqrt 63w ^36 3w^6 sqrt 7 9w^6 sqrt 7 3w^18 sqrt 7 9w^18 sqrt 7

OpenStudy (jhannybean):

\[\large \sqrt{63}*\sqrt{w^{36}}\]

OpenStudy (anonymous):

explain

OpenStudy (anonymous):

\[\sqrt{63w^{36}}= \sqrt{63} * w^{36/2} = w^{18} * \sqrt{9 * 7} = 3w^{18} * \sqrt{7}\]

OpenStudy (jhannybean):

So you have \[\sqrt{63} = \sqrt{3*3*7} \] and \[w ^{36/2}= w^{18}\] combine. \[\large 3\sqrt{7}*w^{18}\]

OpenStudy (jhannybean):

Yeah what Eleven got.....lol

OpenStudy (anonymous):

thank u !! I wish I could give u both a medal lol could u guys help with another ?

OpenStudy (jhannybean):

ok

OpenStudy (anonymous):

what is the exact value of the expression 2 sqrt7 + sqrt14 + 4 sqrt7 ? 4 sqrt7 + sqrt14 6 sqrt7 + sqrt 14 7 sqr28 sqrt28

OpenStudy (jhannybean):

\[2\sqrt{7}+\sqrt{14}+4\sqrt{7}\] just break down the square root. of sqrt 14. \[6\sqrt{7}+\sqrt{14}\]

OpenStudy (jhannybean):

remove tht middle line. lol. that wasnt needed.

OpenStudy (anonymous):

ok thanks what is the exact value of the expression sqrt486 - sqrt24 + sqrt6? simplify is possible 2 sqrt2 2 sqrt6 6 sqrt2 8 sqrt6

OpenStudy (jhannybean):

meant to say you CANT factor out anything from rad 14, so just leave it as it is.

OpenStudy (anonymous):

Oh ok. I'm not too good with sqrt anything.

OpenStudy (jhannybean):

i typed it out wrong -_-

OpenStudy (anonymous):

oh ok. I was wondering where it went.

OpenStudy (jhannybean):

I'm sorry,i'll redo it.Thanks for being patient, heh.

OpenStudy (anonymous):

no problem :)

OpenStudy (jhannybean):

\[\sqrt{486}-\sqrt{24}+\sqrt{6}\]write out the multiples of each first. \[\sqrt{18*27}-\sqrt{6*4}+\sqrt{3*2}\] see if you can pull our any pairs of numbers, if not, simplify further.\[\sqrt{3*3*2}-\sqrt{3*2*2*2}+\sqrt{3*2}\]

OpenStudy (anonymous):

I think its 8 sqrt 6

OpenStudy (anonymous):

right ?

OpenStudy (jhannybean):

sorry, \[\sqrt{18*27}= \sqrt{3*3*2*3*3*3}= 9\sqrt{6}\] ****

OpenStudy (jhannybean):

yes! you are right.

OpenStudy (anonymous):

thank u:)

OpenStudy (jhannybean):

haha np x_x